OpenStudy (anonymous):

Best answer gets a medal! Solve the quadratic equation for it's exact value by using the square root principle 4x^2-13=0

OpenStudy (anonymous):

Do a number line to help ...

OpenStudy (anonymous):

Simplifying 4x2 + -13 = 0 Reorder the terms: -13 + 4x2 = 0 Solving -13 + 4x2 = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '13' to each side of the equation. -13 + 13 + 4x2 = 0 + 13 Combine like terms: -13 + 13 = 0 0 + 4x2 = 0 + 13 4x2 = 0 + 13 Combine like terms: 0 + 13 = 13 4x2 = 13 Divide each side by '4'. x2 = 3.25 Simplifying x2 = 3.25 Take the square root of each side: x = {-1.802775638, 1.802775638}